An object of mass 5 kg travelling from left to right at 6ms-1 collides with an object of mass 4kg which is travelling at 2.5ms-1 from right to left. The 4kg object bounces back at 2ms-1


1
Expert's answer
2021-05-08T11:48:48-0400

"m_1v_1-m_2v_2=m_1v_1'+m_2v_2'\to"


"5\cdot6-4\cdot2.5=5\cdot v_1'+4\cdot2\to v_1'=2.4\ (m/s)" . Answer